winning the jackpot in a particular lottery requires that you select the correct two numbers between 1 and 27 and, in a separate drawing, you must also select the correct single number between 1 and 20. find the probability of winning the jackpot.
Answer by stanbon(75887)

winning the jackpot in a particular lottery requires that you select the correct two numbers between 1 and 27 and, in a separate drawing, you must also select the correct single number between 1 and 20. find the probability of winning the jackpot.
-----------------
# of ways to win: 1*1
# of possible outcomes:: 27C2*20 = 351*20 = 7020
-----
P(win) = 1/7020
=============
Cheers,
Stan H.
---------------